In a 50 mL or 100 mL conical flask, mix 3.0 g (0.022 mol) of salicylic acid and 3.0 mL (3.2 g, 0.031 mol) acetic anhydride; add 6 drops of concentrated sulfuric acid using a clean, dry Pasteur pipette (gloves, safety glasses).

-Heat the solution over a boiling water bath (a large beaker of boiling water on a hot plate and the conical flask is suspended in it) for 10 minutes with occasional swirling; all solids should dissolve during this time. Remove the flask from the water bath and add about 10 g of ice to the hot mixture.

-Swirl for a few minutes to hydrolyse the excess acetic anhydride, and then add 10-15 mL of water. Swirl and cool to complete the precipitation of the crude product.


Q/ Determine the yield of the crude aspirin, show reasoning and full working




1
Expert's answer
2021-10-28T02:55:55-0400

100

3g+3g= 6g

10g+15g=25g

15+6=21g

100-21= 79g

Yield= 79g